The US uses about 8 x 1019 J of energy per year. How much mass would that require if we could totally convert mass to energy? If the mass came from, say, turning used cars entirely into energy, estimate how many would be required?
Answer #1

Energy and Mass are related by,
E = m*c^2
8 * 10^19 = m * (3.0*10^8)^2
m = 888.9 Kg

The average car of mass is around 1000 Kg
So 1 car would do !!

answered by: ANURANJAN SARSAM